# BERP: A Blind Estimator of Room Parameters for Single-Channel Noisy Speech Signals

### v6

@@ III-A1 Intelligibility Parameters
Percentage articulation loss of consonants. considers the proportion of misunderstood consonants, a concept originally introduced by Peutz [9]. Unlike STI, which overlooks listener proficiency and linguistic factors, extends the limitation of STI by not discounting significant information related to intelligibility and considering linguistic proficiency. Thus, emerges as a necessary complement to comprehensively assess speech intelligibility. can be calculated from the STI according to Farrell Becker’s empirical formula [67]:
@@ III-A2 Reverberation Parameters
Reverberation time () and early decay time (EDT) are relevant to reverberation and quantify the subjective impression of the vivacity of a sound field. is the most essential RAP, as it characterizes the physical properties of the RACs for which the reverberation energy is distributed within - dB. The EDT represents the decay time for the initial dB to emphasize the more important contribution of early reflection to perceived reverberation. Both RAPs are derived from the RIR via Schroeder’s back integration method[68]. We use the ISO 3382-1:2009 standard [10] to derive and EDT, keeping with the same approach used in several baselines[33, 37, 38, 39, 29, 35].
@@ III-B2 Sound Source Distance
The sound source distance contributes significantly to complementing the sound source localization (SSL) by integrating it with the DOA of the sound source [41, 40]. SSL is widely used in applications such as sound source separation [1], audio-oriented and navigational systems [6], speech-related applications [73], and human-robot interactions [24]. Furthermore, is intimately related to the perception of speech intelligibility, particularly in terms of [4].
@@ III-C Occupancy Level
Detecting the instantaneous occupancy level around a listener’s location in a room is highly useful for several applications. The audio- and speech-based estimation methods have attracted attention because of their nonintrusiveness, real-time estimation, and cost-effectiveness[46]. The number of occupants affects reverberation [74], thus affecting the efficacy of demand-driven hearing aid systems and speech enhancement methods. In addition, interference speeches generated by the occupants around a listener affect target signals that the listener intends to receive. Knowing the occupancy level helps control interference to achieve intelligible and clear transmission.
In the context of smart homes, the estimated number of nearby occupants can optimize the control of demand-driven heating, ventilation, and air conditioning (HVAC) operations in the local space to significantly reduce the cost of building operations for sustainable smart buildings [25, 26]. In the XR and AR scenarios, the local occupancy level, as an important factor in environmental information factor, is fundamental to ensuring safe interaction in real-world scenes, especially in public spaces populated by others.
@@ IV Preparation of Training Data
This section details the preparation of training and testing data to validate our work. With respect to RAPs and RGPs, we train and evaluate our model using a combination of synthetic and real datasets. For occupancy level estimation, we rely solely on synthetic data because of the lack of real datasets. This section is divided into three subsections. First, we introduce the signal models used to generate synthetic data for blind room parameter estimation. Second, we describe the collections of external databases for generating synthetic and real datasets. Third, we explain how to generate synthetic datasets by using signal models and external databases.
@@ IV-A1 Noisy Reverberant Signal Model
This noisy reverberant signal model is used to generate synthetic data for training the unified module in the proposed framework, as shown in Fig. 2. Details on how this model is used to synthesize the observed reverberant signal are provided in IV-C1. The observed reverberant signal perceived by a listener while transmitting from a speaker within a room and subject to the influence of the background environmental noise, can be formulated as:
@@
IV-A2 Occupancy Reverberant Signal Model
Due to the lack of a reverberant speech corpus for blind occupancy level estimation, we propose a signal model to simulate nearby speakers around a listener to generate synthetic data to train the occupancy module in the proposed framework (Fig. 2) to estimate the instantaneous occupancy level around the listener. Details of the data generation pipeline are given in Section IV-C2.
@@ IV-A2 Occupancy Reverberant Signal Model
where signifies the occupancy reverberant speech signal, represents the speech signal originating from the -th speaker proximal to the listener, and denotes the distance between the -th speaker and the listener, which adheres to a Gaussian distribution. represents the total count of occupants around an observer’s location. represents the baseline amplitude observed at a distance of from the listener, i.e., the original amplitude of the speech without attenuation. Here, is assumed to be equal to , which is a proper choice considering the preferred interpersonal distance[75]. denotes the background noise. denotes the RIR of each occupant with amplitude normalization. The proposed model assumes that the distance from the -th speaker to the listener should exceed , (i.e., it defines a minimal circular boundary around the speaker based on the preferred interpersonal distance[75]), and the amplitude corresponding to each occupant decays proportionally according to the multiplicative inverse of the associated distance[25].
Currently, we focus on the overall effect of reverberation on overlapping speech, and given the difficulty of collecting real RIRs for every occupant, we treat early reflections as part of the direct speech signal, as early reflections can be integrated as direct sounds under noisy conditions [76, 77, 78]. In addition, McKenzie et al.[79] demonstrated that positional information is not reflected primarily in late reverberation. Therefore, the signal model can be simplified by integrating early reflections of into speech signals and approximating as a single global RIR . Then, Eq. (6) is simplified as:
We agree that incorporating early reflections with full position dependence would allow for a more accurate estimation of real-world scenarios. A model that explicitly accounts for early reflection will be addressed in future work.
@@ IV-B Dataset Collection
This subsection details the collection of external datasets for the next step of dataset compilation. According to Eqs. (5) and (7), three datasets are required to synthesize noisy reverberant and occupancy signals: RIRs, clean speech recordings, and background noise datasets. For blind estimation of RAPs and RGPs, these datasets were used to generate synthetic data (noisy reverberant speech signals) following Eq. (5). In addition, real reverberant speech recordings were collected to improve the practical adaptability of the model. Synthetic and real datasets were combined to train the unified module and evaluated separately on synthetic and real data. For the occupancy level estimation task, due to the lack of real recorded occupancy data, the occupancy module was trained and evaluated solely on synthetic data generated via Eq. (7).
@@ IV-B1 RIR Dataset Collection
We collected real RIRs from five existing real RIR datasets to synthesize observed speech signals, including the noisy reverberant signal and the occupancy reverberant signal . These RIR datasets include the Arni dataset [80], the Motus dataset [81], the BUT ReverbDB [82], the ACE corpus [32], and the OpenAIR [83]. Each dataset comprises monochanneled and omnidirectionally recorded RIRs. All the selected datasets contain the detailed attributes in terms of room volume, including 39 rooms with volumes ranging from 30 to , locations of sound sources and receivers or equivalent information such as sound source distances. These RIRs cover a wide range of acoustic environments, featuring various volumes and geometries of the room, varied locations of the sound source and receiver, and different sound absorption coefficients of the room surfaces. Hence, it can contain a wide spectrum of broadband RAPs and RGPs. We resampled all RIRs to kHz.
@@ IV-B4 Real Reverberant Speech Recordings Collection
For the blind estimation task of RAPs and RGPs, we mixed a smaller amount of real reverberant speech recordings with large-scale synthetic reverberant data to train and evaluate the model. In total, we collected 11,160 samples from the BUT Retransmission dataset [82] that transmit clean speech in real rooms to obtain real recordings.
@@ IV-C1 Synthesis Pipeline of Noisy Reverberant Speech Signals
In the collected real RIR dataset with labels of RAPs and RGPs, we further employ a data augmentation strategy. The strategy involves data upsampling and downsampling techniques to modulate the distribution of labels. For data upsampling, we repeat some RIRs in which those labels appear less frequently to convolve with more clean speech recordings. Because each clean speech recording is distinct, we still guarantee that generated examples have enough diversity after data upsampling. Conversely, for data downsampling, we downsample those RIRs (skipping someone) with abundant labels. The degrees of upsampling and downsampling are calibrated based on the relative rarity of labels. For example, collected real RIR datasets often include many labeled examples with smaller volumes (e.g., ) but very few with larger volumes (e.g., ), resulting in an imbalanced distribution of room volume labels. To address this imbalance, we downsampled RIRs associated with more prevalent room volume labels while concurrently upsampling RIRs corresponding to the rarer volume labels. This data augmentation strategy is open-sourced in the data preprocessing pipeline333https://github.com/Alizeded/BERP/blob/main/notebooks/dataset_preprocess.ipynb. Afterward, a comprehensive collection of real RIRs is successfully compiled. This RIR dataset contains a wide range of RIRs, for which the corresponding spans from to s. The mean reverberation time () across this dataset is approximately s.
@@
We randomly sample 47,430 clips from the LibriSpeech corpus as clean speech signals, regardless of the speaker information and linguistic content they contain. In parallel, noise signals are randomly sampled, following an independent and identically distributed (I.I.D.) pattern, from the DEMAND and BUT datasets. Then, according to Eq. (5), we synthesize the noisy reverberant speech signals via RIR, clean speech, and noise signals, for a total of 261 hours. To enhance the robustness and efficacy of the model in various noisy environments, the signal-to-noise ratio (SNR) between reverberant and noisy signals is uniformly varied by adjusting the SNR at five different levels, ranging from 0 to 20dB in 5dB increments, including a scenario without noise (Inf). Given the uniqueness of each clip, we guarantee that every synthesized speech signal maintains its individuality in terms of both its waveform and linguistic content, further augmenting the diversity and richness of the synthetic dataset.
IV-C2 Synthesis Pipeline of Occupancy Reverberant Speech Signals
@@
First, we applied voice activity detection to the LibriSpeech corpus to segment and annotate timestamps corresponding to speech and silence segments. This process underlies annotations of the synthesized occupancy reverberant signals.
@@
Next, we model each occupant’s distance from the listener via a Gamma distribution , where and are the shape and scale parameters, respectively. To validate this setting, we fit real occupancy-distance data from [86] (see Fig. 3) and used a chi-square goodness-of-fit test, resulting in a p-value above 0.05. This finding indicates that a Gamma distribution effectively models the occupancy distance. Consequently, we set and . Furthermore, we limit the occupancy distance up to 6m, according to the considerations discussed below. The normal speech level at 1m (56–58dB[87]) attenuates to 34–36dB at 6m (5m distance)444Calculated based on ISO 9613-2:2024[88] without considering reflections in a typical room setting at 23 and humidity [89].. In addition, given the early reflection effect for speech intelligibility with a 4dB improvement at a 5m distance in noise conditions[78], a speech sound level of 6m places at 38–42dB. Such levels are easily masked by background noise in most daily scenarios, so we treat occupant speech beyond this distance as part of background noise.
@@
At the same time, we used statistics on the real-world occupancy level and employed kernel density estimation (KDE) [90] to determine the occupants (i.e., the number of individuals) for each synthesized overlapping sample. Note that this number represents the speech numbers used to synthesize overlapping speech rather than the instantaneous occupancy level at each time frame. After analyzing a large-scale real occupancy level statistics from [91], we determined that traditional parametric distributions (e.g., discretized Gamma or negative binomial) do not adequately capture complexity of the observed patterns. Therefore, we instead use a nonparametric approach to model the occupancy level distribution. The chi-square test (p-value0.05) confirms that KDE effectively represents the underlying distribution (see Fig. 4). Then, we use this validated KDE model to generate occupancy levels that closely mirror real-world settings. Figure 4 depicts the resulting occupancy distribution in the synthetic dataset compiled. Here, we set the maximum to 17 within the 6m (max ) radius, based on the rationale presented below. First, given that the area of this radius is approximately 110.0 and that the preferred social distance is m (6.2) per person [75], 17 is a reasonable capacity in this area. Second, given the difficulty of detecting occupancy levels from overlapping speech, 17 is the maximum occupancy level that we can now detect. Notably, these considerations are a plausible practical guide rather than a definitive scientific conclusion. Real environments may exceed this threshold, and more studies are needed to validate them under various acoustic conditions.
@@
Finally, via Eq. (7), we synthesize occupancy reverberant speech by overlapping speech signals uniformly sampled from the LibriSpeech corpus, convolving with real RIRs, adding background noise, and aligning with annotated speech and silence segments. The background noise is randomly sampled from the compiled noise dataset mentioned in Section IV-B3. The noise SNR is uniformly adjusted in six levels from dB to dB in the dB step, to align with of that background noise originates away from the 6-meter circular area. The initiation index for each overlapping speech signal is determined based on an I.I.D. pattern. Additionally, to replicate the local room acoustics, the room volumes in the RIRs are precisely matched with their corresponding RIRs, thereby ensuring a realistic acoustic environment. Furthermore, when convolving the RIRs, we preprocessed the RIRs to discard position-dependent information. Specifically, we filtered the direct sound and early reflection (initial ms) of the RIR. We also normalize the amplitude of to eliminate the cues related to the sound source distances from the RIR amplitudes. These steps discard position-dependent information to the greatest extent feasible, ensuring that the RIR used in this signal model represents only late reverberation-related acoustical characteristics. We obtain a dataset comprising 47,430 samples of occupancy reverberant speech signals, ranging from 10 to 25s, for a total of 222 hours of audio. Figure 5 shows an example of an occupancy reverberant speech signal and its corresponding instantaneous occupancy level.

### v7

@@ III-A1 Intelligibility Parameters
Percentage articulation loss of consonants. considers the proportion of misunderstood consonants, a concept originally introduced by Peutz [9]. Unlike the STI, which overlooks listener proficiency and linguistic factors, extends the limitation of STI by not discounting significant information related to intelligibility and considering linguistic proficiency. Thus, emerges as a necessary complement to assess speech intelligibility comprehensively. can be calculated from the STI according to Farrell Becker’s empirical formula [68]:
@@ III-A2 Reverberation Parameters
Reverberation time () and early decay time (EDT) are relevant to reverberation and quantify the subjective impression of the vivacity of a sound field. is the most essential RAP, as it characterizes the physical properties of the RACs for which the reverberation energy is distributed within - dB. The EDT represents the decay time for the initial dB to emphasize the more important contribution of early reflection to perceived reverberation. Both RAPs are derived from the RIR via Schroeder’s back integration method[69]. We use the ISO 3382-1:2009 standard [10] to derive and EDT, following the same approach adopted in several baseline studies[34, 38, 39, 40, 29, 36].
@@ III-B2 Sound Source Distance
The sound source distance significantly contributes to complementing the sound source localization (SSL) by integrating it with the DOA of the sound source [42, 41]. SSL is widely used in applications such as sound source separation [1], audio-oriented and navigational systems [6], speech-related applications [74], and human-robot interactions [24]. Furthermore, is intimately related to the perception of speech intelligibility, particularly in terms of [4].
@@ III-C Occupancy Level
Detecting the instantaneous occupancy level around a listener’s location in a room is highly useful for several applications. Audio- and speech-based estimation methods have attracted attention because of their nonintrusiveness, real-time estimation, and cost-effectiveness[47]. The number of occupants affects reverberation [75], thus affecting the efficacy of demand-driven hearing aid systems and speech enhancement methods. In addition, interference speeches generated by the occupants around a listener affect target signals that the listener intends to receive. Knowing the occupancy level helps control interference to achieve intelligible and clear transmission.
In the context of smart homes, the estimated number of nearby occupants can optimize the control of demand-driven heating, ventilation, and air conditioning (HVAC) operations in the local space to significantly reduce the cost of building operations for sustainable smart buildings [25, 26]. In the XR and AR scenarios, the local occupancy level, as a key component of environmental information factor, is fundamental to ensuring safe interaction in real-world scenes, especially in public spaces populated by others.
@@ IV Preparation of Training Data
This section details the preparation of training and testing data to validate our work. For the task of RAPs and RGPs estimation, we train and evaluate the model using a combination of synthetic and real datasets. For the occupancy level estimation task, we rely solely on synthetic data because of the lack of real datasets. This section has three subsections. First, we introduce the signal models used to generate synthetic data for these two tasks. Second, we describe the collections of external databases used for generating synthetic and real datasets. Third, we explain how to generate synthetic datasets for the two tasks using signal models and external databases.
@@ IV-A1 Noisy Reverberant Signal Model
The noisy reverberant signal model is used to generate synthetic data for training the unified module in the proposed method, as shown in Fig. 2. Details on how this model is used to synthesize the observed reverberant signal are provided in IV-C1. The observed reverberant signal perceived by a listener while transmitting from a speaker within a room and subject to the influence of the background environmental noise, can be formulated as:
@@
IV-A2 Multitalker Reverberant Signal Model
Due to the lack of a reverberant speech corpus for blind occupancy level estimation, we propose a signal model to simulate nearby speakers around a listener to generate synthetic data to train the occupancy module in the proposed method (Fig. 2) to estimate the instantaneous occupancy level around the listener. Details of the data generation pipeline are given in Section IV-C2.
@@ IV-A2 Multitalker Reverberant Signal Model
where signifies the multitalker reverberant speech signal, represents the speech signal originating from the -th speaker proximal to the listener, and denotes the distance between the -th speaker and the listener, which adheres to a Gaussian distribution. represents the total count of occupants around an observer’s location. represents the baseline amplitude observed at a distance of from the listener, i.e., the original amplitude of the speech without attenuation. Here, is assumed to be equal to , which is a proper choice considering the preferred interpersonal distance[76]. denotes the background noise. denotes the RIR of each occupant with amplitude normalization. The proposed model assumes that the distance from the -th speaker to the listener should exceed , (i.e., it defines a minimal circular boundary around the speaker on the basis of the preferred interpersonal distance[76]), and the amplitude corresponding to each occupant decays proportionally according to the multiplicative inverse of the associated distance[25].
Currently, we focus on the overall effect of reverberation on overlapping speech. Given the difficulty of collecting real RIRs for every occupant, we treat early reflections as part of the direct speech signal, as they can be integrated with it under noisy conditions [77, 78, 79]. In addition, McKenzie et al.[80] demonstrated that positional information is not primarily conveyed in late reverberation. Therefore, the signal model can be simplified by integrating early reflections of into speech signals and approximating as a single global RIR . Then, Eq. (6) is simplified as:
We acknowledge that incorporating early reflections with full position dependence would allow for a more accurate estimation of real-world scenarios. A model that explicitly accounts for early reflection will be addressed in future work.
@@ IV-B Dataset Collection
This subsection details the collection of external datasets for the next step of dataset compilation. According to Eqs. (5) and (7), three datasets are required to synthesize noisy reverberant and occupancy signals: RIRs, clean speech recordings, and background noise datasets. For blind estimation of RAPs and RGPs, these datasets were used to generate synthetic data (noisy reverberant speech signals) following Eq. (5). In addition, real reverberant speech recordings were collected to improve the practical adaptability of the model. Synthetic and real datasets were combined to train the unified module, which was then evaluated separately on synthetic and real data. For the occupancy level estimation task, due to the lack of real recorded occupancy data, the occupancy module was trained and evaluated solely on synthetic data generated via Eq. (7).
@@ IV-B1 RIR Dataset Collection
We collected real RIRs from five existing real RIR datasets to synthesize observed speech signals, including the noisy reverberant signal and the occupancy reverberant signal . These RIR datasets include the Arni dataset [81], the Motus dataset [82], the BUT ReverbDB [83], the ACE corpus [32], and the OpenAIR [84]. Each dataset comprises monochanneled and omnidirectionally recorded RIRs. All the selected datasets contain the detailed attributes in terms of room volume, covering 39 rooms with volumes ranging from 30 to , locations of sound sources and receivers or equivalent information such as sound source distances. These RIRs cover a wide range of acoustic environments, featuring various volumes and geometries of the room, varied locations of the sound source and receiver, and different sound absorption coefficients of the room surfaces. Hence, they can contain a wide spectrum of broadband RAPs and RGPs. We annotated all the labels for the collected RIRs using calculation methods in Section III. While labels were derived via the empirical Eq. (1) due to a lack of direct measurement, its limited applicability motivated us to predict end-to-end, aiming for adaptability to potential future scenarios where might be obtained directly, not just via the STI. We resampled all RIRs to kHz.
@@ IV-B4 Real Reverberant Speech Recordings Collection
For the blind estimation task of RAPs and RGPs, we mixed a smaller number of real reverberant speech recordings with large-scale synthetic reverberant data to train and evaluate the model. In total, we collected 11,160 samples from the BUT Retransmission dataset [83], where clean speech is transmitted in real rooms to obtain real recordings.
@@ IV-C1 Synthesis Pipeline of Noisy Reverberant Speech Signals
In the collected real RIR dataset with labels of RAPs and RGPs, we further employ a data augmentation strategy. The strategy involves data upsampling and downsampling techniques to modulate the distribution of labels. For data upsampling, we repeat some RIRs in which those labels appear less frequently to convolve with more clean speech recordings. Because each clean speech recording is distinct, we still guarantee that the generated examples have enough diversity after data upsampling. Conversely, for data downsampling, we skip some RIRs that have abundant labels. The degrees of upsampling and downsampling are calibrated based on the relative rarity of labels. For example, collected real RIR datasets often include many labeled examples with smaller volumes (e.g., ) but very few with larger volumes (e.g., ), resulting in an imbalanced distribution of room volume labels. To address this imbalance, we downsampled RIRs associated with more prevalent room volume labels while concurrently upsampling RIRs corresponding to the rarer volume labels. This data augmentation strategy is open-sourced in the data preprocessing pipeline333https://github.com/Alizeded/BERP/blob/main/notebooks/dataset_preprocess.ipynb. Afterward, a comprehensive collection of real RIRs is successfully compiled. This RIR dataset contains a wide range of RIRs, for which the corresponding spans from to s. The mean reverberation time () across this dataset is approximately s.
@@
We randomly sample 47,430 clips from the LibriSpeech corpus as clean speech signals, regardless of the speaker information and linguistic content they contain. In parallel, noise signals are randomly sampled, following an independent and identically distributed (I.I.D.) pattern, from the DEMAND and BUT datasets. Then, according to Eq. (5), we synthesize the noisy reverberant speech signals via RIR, clean speech, and noise signals, for a total of 261 hours. To increase the robustness and efficacy of the model in various noisy environments, the signal-to-noise ratio (SNR) between reverberant and noisy signals is uniformly varied by adjusting the SNR at five different levels, ranging from 0 to 20dB in 5dB increments, including a scenario without noise (Inf). Given the uniqueness of each clip, we guarantee that every synthesized speech signal maintains its individuality in terms of both its waveform and linguistic content, further augmenting the diversity and richness of the synthetic dataset.
IV-C2 Synthesis Pipeline of Multitalker Reverberant Speech Signals
@@
First, we applied voice activity detection to the LibriSpeech corpus to segment and annotate timestamps corresponding to speech and silence segments. This process underlies annotations of the synthesized multitalker signals.
@@
Next, we model each occupant’s distance from the listener via a Gamma distribution , where and are the shape and scale parameters, respectively. To validate this setting, we fit real occupancy-distance data from [87] (see Fig. 3) and used a chi-square goodness-of-fit test, resulting in a p-value above 0.05. This finding indicates that a Gamma distribution effectively models the occupancy distance. Consequently, we set and . Furthermore, we limit the occupancy distance up to 6m, according to the considerations discussed below. The normal speech level at 1m (56–58dB[88]) attenuates to 34–36dB at 6m (5m distance)444Calculated based on ISO 9613-2:2024[89] without considering reflections in a typical room setting at 23 and humidity [90].. In addition, given the early reflection effect for speech intelligibility with a 4dB improvement at a distance of 5m in noisy conditions[79], the speech sound level at 6m is 38–42dB. Such levels are easily masked by background noise in most daily scenarios, so we treat occupant speech beyond this distance as part of background noise.
@@
Meanwhile, we used statistics on the real-world occupancy level and employed kernel density estimation (KDE) [91] to determine the occupants (i.e., the number of individuals) for each synthesized overlapping sample. Note that this number represents the speech numbers used to synthesize overlapping speech rather than the instantaneous occupancy level at each time frame. After analyzing a large-scale real occupancy level statistics from [92], we determined that traditional parametric distributions (e.g., discretized Gamma or negative binomial) do not adequately capture the complexity of the observed patterns. Therefore, we instead use a nonparametric approach to model the occupancy level distribution. The chi-square test (p-value0.05) confirms that KDE effectively represents the underlying distribution (see Fig. 4). Then, we use this validated KDE model to generate occupancy levels that closely mirror real-world settings. Figure 4 depicts the resulting occupancy distribution in the compiled synthetic dataset. Here, we set the maximum to 17 within the 6m (max ) radius, on the basis of the rationale presented below. First, given that the area of this radius is approximately 110.0 and that the preferred social distance is m (6.2) per person [76], 17 is a reasonable capacity in this area. Second, given the difficulty of detecting occupancy levels from overlapping speech, 17 is the maximum occupancy level that we can now detect. Notably, these considerations are a plausible practical guide rather than a definitive scientific conclusion. Real environments may exceed this threshold, and more studies are needed to validate these findings under various acoustic conditions.
@@
Finally, via Eq. (7), we synthesize occupancy reverberant speech by overlapping speech signals uniformly sampled from the LibriSpeech corpus, convolving with real RIRs, adding background noise, and aligning with annotated speech and silence segments. The background noise is randomly sampled from the compiled noise dataset mentioned in Section IV-B3. The noise SNR is uniformly adjusted across six levels, ranging from dB to dB in the dB increments, to align with of that background noise originates away from the 6-meter circular area. The initiation index for each overlapping speech signal is determined on the basis of an I.I.D. pattern. Additionally, to replicate the local room acoustics, the room volumes in the RIRs are precisely matched with their corresponding RIRs, thereby ensuring a realistic acoustic environment. Furthermore, when convolving the RIRs, we preprocessed the RIRs to discard position-dependent information. Specifically, we filtered the direct sound and early reflection (initial ms) of the RIR (see Section IV-A2 for more details). We also normalize the amplitude of to eliminate the cues related to the sound source distances from the RIR amplitudes. These steps discard position-dependent information to the greatest extent feasible, ensuring that the RIR used in this signal model represents only late reverberation-related acoustical characteristics. We obtain a dataset comprising 47,430 samples of multitalker reverberant speech signals, ranging from 10 to 25s, for a total of 222 hours of audio. Figure 5 shows an example of a multitalker reverberant speech signal and its corresponding instantaneous occupancy level.